OpenStudy (anonymous):

find the equation for the line satisfying the given condition. Perpendicular to 5x+2y=3 and passing through (-4,-1)

OpenStudy (raden):

the equation of straight line which perpendicular with line ax + by = c and passes through the point (x1,y1) is bx - ay = b(x1) - a(y1) so, the equation of line which perpendicular with line 5x + 2y = 3 and passes through the point (-4,-1) is 2x - 5y = 2(-4) - 5(-1) or it can be simplied to 2x - 5y = -3

OpenStudy (raden):

does that helps ?

OpenStudy (anonymous):

yes thnx

OpenStudy (raden):

you're welcome